[Brahms-l] Shift report 20050507 08:00-16:00

From: <apache@brahms-web.brahms.bnl.gov>
Date: Sat May 07 2005 - 15:59:36 EDT
Shifter leader was rw
Shifters were 


1. Beam for 6 hours. 
   From the previous shift to 8:30, 
   and 10:30 to the end of this shift.
   
2. Data run 14486 to 14496. No pedestal runs.

3. No detector problem.

_______________________________________________
Brahms-l mailing list
Brahms-l@lists.bnl.gov
http://lists.bnl.gov/mailman/listinfo/brahms-l
Received on Sat May 7 15:59:48 2005

This archive was generated by hypermail 2.1.8 : Sat May 07 2005 - 16:00:00 EDT